Find the area between the curve y=1/x and y=1/x+x^2 from x=2 to x=2.


1
Expert's answer
2021-04-14T14:55:36-0400

The area between the curves "y=\dfrac{1}{x}" and "y=x^{2}+\dfrac{1}{x}" is from x=a to x=b is


"\int_{a}^{b} (x^{2}+\dfrac{1}{x}) dx-\int_{a}^{b}\dfrac{1}{x}dx" ="\int _{a}^{b} x^{2} dx=[\dfrac{x^{3}}{3}]_{a}^{b}=\dfrac{b^{3}}{3}-\dfrac{a^{3}}{3}"


From x=2 to x=2 the area is clearly zero.